OpenStudy (anonymous):

Find the exact value by using a half-angle identity. tangent of seven pi divided by eight.

OpenStudy (danjs):

first write the half- angle identities down

OpenStudy (danjs):

and recall tan(x) = sin(x) / cos(x)

OpenStudy (anonymous):

finally some help.

OpenStudy (danjs):

finally a question that is not solve y=mx+b or simple arithmatic

OpenStudy (anonymous):

Lol

OpenStudy (danjs):

\[\cos^2(u) = \frac{ 1+\cos(2u) }{ 2 }\] and \[\sin^2(u) = \frac{ 1-\cos(2x) }{ 2 }\]

OpenStudy (anonymous):

Ok, give me a min.

OpenStudy (danjs):

so

OpenStudy (danjs):

\[\tan^2(u) = \frac{ \sin^2(u) }{ \cos^2(u) }\]

OpenStudy (jhannybean):

I would just divide the two identities.

OpenStudy (danjs):

u = 7pi/8 in this problem

OpenStudy (danjs):

what you have so far?

OpenStudy (anonymous):

Find all solutions to the equation. sin2x + sin x = 0

OpenStudy (anonymous):

wait

OpenStudy (danjs):

\[\tan^2(u) = \frac{ 1-\cos(2u) }{ 1+\cos(2u) }\]

OpenStudy (anonymous):

Hold on.

OpenStudy (danjs):

\[\tan(u) = \sqrt{\frac{ 1-\cos(2u) }{ 1+\cos(2u) }}\]

OpenStudy (danjs):

k

OpenStudy (danjs):

recall cos(2u) = cos(2*7pi/8) = cos(7pi/4) = \[\frac{ \sqrt{2} }{ 2 }\]

OpenStudy (danjs):

\[\sqrt{\frac{ 1-\frac{ \sqrt{2} }{ 2 } }{ 1+\frac{ \sqrt{2} }{ 2 } }}\]

OpenStudy (jhannybean):

\[\tan(x) = \frac{\sin(x)}{\cos(x)} \implies \tan^2(x) = \frac{\sin^2(x)}{\cos^2(x)}\]\[sin^2(x) = \frac{1}{2}(1-\cos(2x))\]\[\cos^2(x) = \frac{1}{2}(1+cos(2x))\]\[\large\tan^2(x)= \frac{\frac{1}{2}(1-\cos(2x))}{\frac{1}{2}(1+\cos(2x))}\]

OpenStudy (danjs):

yeah i just did that

OpenStudy (jhannybean):

That's what I was trying to type up, haha,

OpenStudy (danjs):

it is the negative root of the last thing i typed

OpenStudy (anonymous):

ok wait, I get it. But im still confused.

OpenStudy (danjs):

which part?

OpenStudy (jhannybean):

\[\large -\sqrt{\frac{ 1-\frac{ \sqrt{2} }{ 2 } }{ 1+\frac{ \sqrt{2} }{ 2 } }}\]

OpenStudy (anonymous):

ok well... Dude, I really dislike this class.

OpenStudy (danjs):

yes

OpenStudy (danjs):

yeah it is correct, both tan 7pi/4 and the root we found is approx -0.414

OpenStudy (danjs):

it is just screwin around with the 5 or 10 identities you learn

OpenStudy (danjs):

no

OpenStudy (anonymous):

give me a min, to solve.

OpenStudy (anonymous):

Ok I give up, I'm have to really study this some other time. Sorry